# Certified Randomness implies Secure Classical Position-Verification

### v2

@@ Appendix B Instantiation
A quantum algorithm $\mathcal{A}$ given $C\sim\mathcal{D}$ is said to solve $b$-$\mathsf{XHOG}$ if it outputs a bitstring $z$ such that
$\underset{C\sim\mathcal{D}}{\mathbb{E}}\left[\underset{z\sim\mathcal{A}^{C}}{\mathbb{E}}\left[p_{C}(z)\right]\right]\geq\frac{b}{N}.$
@@
Specifically, for the single round analysis, RCS-based certified randomness yields different bounds on the von Neumann entropy under different models. Denote the single round output as $A$, the challenge as $C$, and any side information as $E$.
One model that is considered is the semi-honest adversary, where the prover may be entangled with an guesser but performs ideal quantum measurement.
Theorem 12 (Theorem 6.4 of [AH23]).
Consider a semi-honest adversary performing ideal measurement on a state sharing some entanglement with register $E$ while solving $b$-$\mathsf{XHOG}$ for $b\geq\frac{(1-\varepsilon)N}{N+1}$, $A$ is a length-$n$ bitstring, and $C$ is an $n$-qubit quantum circuit from the Haar measure. We have
$\underset{C\sim\mathrm{Haar}(N)}{\Pr}\left[H(A|E)_{\psi}\geq(0.99-\varepsilon)n\right]\geq 1-O(N^{-0.02}).$ (87)
where $\psi$ is the output state.
Finally, a fully general device given oracle access to $C$ is considered.
@@
B.1 Protocols
The above single round results allow them to use the entropy accumulation theorem to define a multi-round protocols that outputs certified smooth min-entropy for each case. Moreover, for randomness expansion, fresh randomness is only consumed on logarithmically many rounds to generate fresh challenge circuits.
We see that the semi-honest adversary model and the random oracle model allows one to obtain a bound on the von Neumann entropy with quantum side information, and the respective RCS-based certified randomness protocols are certified randomness from entropy accumulation protocols satisfying Definition 7. We have omitted discussions of a general device in Section 5 of [AH23] since does not consider quantum side information, and Definition 7 is not satisfied.
We describe below a protocol in Fig. 9 that instantiates CVPV based on Theorem 12, which is closely related to Fig. 2 of [AH23]. The completeness of the protocol follows from the completeness of the original protocol in [AH23], and the soundness follows from the application of the entropy accumulation theorem.
We note that [AH23] claims to prove entropy accumulation of the protocol in Theorem 6.6 from Corollary 6.5, which is incorrect. Since the theorem states entropy is $\Omega(mn)$ with overwhelming probability over choices of $C$, where $m$ is the number of rounds, the proof should be built on Theorem 6.4 (or Theorem 12 in reproduced here). It is straightforward to apply Corollary 6.5 for entropy accumulation, however, but the resulting theorem should no longer contain the clause ‘with probability $1-2^{-\Omega(n)}$ over the choices of $C$’ and the entropy needs to be conditioned on the challenge circuits as well.
  Input: the qubit count $n$, the number of rounds $\ell$, the score parameter $\delta\in[0,1]$ and the fraction of test rounds $\gamma=O((\log n)/\ell)$.
@@ B.1 Protocols
$V$ sample $C\sim\mathrm{Haar}(N)$ and sends $C$ to $P$.
@@
  1. 1.
    $V_{1}$ receives $A_{i}\in\{0,1\}^{n}$ and $V_{2}$ receives $A_{i}^{\prime}\in\{0,1\}^{n}$
  2. 2.
    $V$ aborts if the timing requirement fails
@@
    $V$ aborts if $A_{i}\neq A_{i}^{\prime}$
$V$ samples $T_{i}\sim\mathsf{Bernoulli}(\gamma)$ for all $i\in[\ell]$.
Let $t=|\{i:T_{i}=1\}|$. $V$ aborts if $\frac{1}{t}\sum_{i:T_{i}=1}p_{C}(A_{i})<(1+\delta)/N$.
Figure 9: CVPV protocol against a semi-honest adversary.
For an instantiation based on Theorem 13, we use following protocol in Fig. 10. However, the protocol in Fig. 10 is very different from the protocol in Fig. 4 of [AH23]. Specifically, we do not reuse the circuit and therefore do not perform the test by summing the scores over epochs. The main reason for this choice is that we do not believe the application of entropy accumulation in [AH23] is correct, which we discuss in the supplement Section B.2.
Input: the qubit count $n$, the number of rounds $\ell$, the score parameter $\delta\in[0,1]$ and the fraction of test rounds $\gamma=O((\log n)/\ell)$.
Protocol:
For $i\in[\ell]$:
  1. 1.
    $V$ sample $C_{i}\sim\mathrm{Haar}(N)$ and sends $C_{i}$ to $P$.
  2. 2.
    $V_{1}$ receives $A_{i}\in\{0,1\}^{n}$ and $V_{2}$ receives $A_{i}^{\prime}\in\{0,1\}^{n}$
  3. 3.
    $V$ aborts if the timing requirement fails
  4. 4.
    $V$ aborts if $A_{i}\neq A_{i}^{\prime}$
$V$ samples $T_{i}\sim\mathsf{Bernoulli}(\gamma)$ for all $i\in[\ell]$.
Let $t=|\{i:T_{i}=1\}|$. $V$ aborts if $\frac{1}{\ell}\sum_{i:T_{i}=1}p_{C_{i}}(A_{i})<(1+\delta)/N$.
@@
In both cases, the original entropy accumulation theorem does not apply since Theorem 12 and 13 give entropy lower bounds for some scores, and the score is averaged over all rounds. This is in contract to bounding the entropy given some probability distribution. We discuss how to address this issue in the supplement Section B.3.
B.2 Issues with the Protocol in [AH23]
@@
One may argue that this type of argument could be used against certified randomness based on post-quantum secure trapdoor claw-free functions [BCM+21b]. Indeed, [BCM+21b] explicitly discusses this issue that entropy accumulation theorem requires that single-round entropy bound for all possible input states, including those that are computationally inefficient strings. As a result, [BCM+21b] presents significant additional analysis to show that entropy accumulates in the protocol, and nontrivial work was presented in [MAF23] to use the entropy accumulation theorem.
@@
B.3 Issues with Entropy Accumulation
The single-round entropy lower bounds are in Theorem 12 and 13 are conditioned on the output achieving some score (e.g. $b$ in $b$-$\mathsf{XHOG}$), and the min-tradeoff functions used in [AH23] are defined for continuous scores. As a result, [AH23] developed a modified entropy accumulation theorem, which requires the Markov chain condition. However, for the arguments in Section A, we prove that the modified adversary satisfies the non-signalling condition which allows us to apply generalized entropy accumulation. To apply the entropy accumulation theorem in [AH23] to the modified adversary, we have to show that the modified adversary also satisfies the Markov chain condition.
@@ B.3 Issues with Entropy Accumulation
For the quantum channel defined in E.q. 90, for some $\mathcal{R}^{\prime}$, the Markov chain condition $A_{i}\leftrightarrow I_{i}\leftrightarrow I_{i+1}$ is satisfied:
@@
As a result, we can bound $H_{\rm min}^{\varepsilon}(A_{1}^{n}|C_{1}^{n}G_{1}^{n})$ by bounding $H_{\rm min}^{\varepsilon}(A_{1}^{n}|I_{1}^{n})$ instead with the entropy accumulation theorem of [AH23]. We can expand the definition of certified randomness from entropy accumulation to the case where the min-tradeoff function is for a score. To do this, we simply need to change $q$ for $\Sigma_{i}(q)$ into a score $s$, and define $h$ as $h=\min_{s}f(s)$. With this expanded definition, the protocols in Fig. 9 and 10 are certified randomness from repetition due to Theorem 12 and 13, and they are therefore CVPV protocols.
@@
Appendix C Missing Proofs
In this section, we provide the additional proofs missing from the main body for brevity and due to similarity with already provided proofs.
C.1 Proof of Theorem 8
@@
As in the proof of Theorem 7, we will consider hybrid experiments:
  • Hybrid 0: This is the original CVPV soundness experiment.
    with $u_{i}\leftarrow\{0,1\}^{n}$ being a fresh random string.
Claim 11.
$p_{1}\geq p_{0}$.
Proof.
Claim 12.
$p_{2}\geq p_{1}$.
Proof.
Claim 13.
Proof.
  • $\widetilde{\mathcal{B}}$ simulates $\mathcal{B}$.
Claim 14.
$p_{3}\geq p_{2.\ell}$.
Proof.
Claim 15.
$p_{3}\leq\mathsf{negl}(\lambda).$
Proof.
Follows from the certified randomness property (Definition 3). $\mathcal{A}$ plays the role of the prover and $\mathcal{B}$ that of the eavesdropper. Note that the no-signalling condition is satisfied because $(\ell-1)\Delta<1$, which means no message can be sent between $\mathcal{A}$ and $\mathcal{B}$ in time from the first challenge until the last. ∎

### v3

@@ 8.1 Instantiation Using Random circuit sampling
A quantum algorithm $\mathcal{A}$ given $C\sim\mathcal{D}$, where $\mathcal{D}$ is some distribution of $\log_{2}N$-qubit quantum circuits, is said to solve $b$-$\mathsf{XHOG}$ if it outputs a bitstring $z$ such that
$\underset{C\sim\mathcal{D}}{\mathbb{E}}\left[\underset{z\sim\mathcal{A}^{C}}{\mathbb{E}}\left[p_{C}(z)\right]\right]\geq\frac{b}{N},$
where $p_{C}(z)=|\langle 0|C|z\rangle|^{2}$.
@@
However, since the test condition requires passing $\mathsf{XHOG}$, exponentially costly classical computation of the probability amplitudes of the received bitstrings is required. As a result, a drawback of this protocol is the asymmetry in the computational power of the adversary and the verifier, namely the adversary is polynomial-time and the verifier can perform exponentially expensive verification.
The polynomial-time constraint on the verifier applies explicitly for the analysis in Section 5 of [AH23] where the Long List Hardness Assumption, a new computational hardness assumption, is used. We do not base our analysis on this model due to reasons we will discuss below. The polynomial-time constraint is more subtly incorporated in the analysis in Section 7 of [AH23], where the adversary is given oracle access to the challenge circuits. This model aims to capture the situation where the explicit description of the circuit cannot help the adversary in any way other than allowing the direct execution of the circuit, which also precludes the possibility of computing probability amplitudes. This is the adversary model we consider for our instantiation of the CVPV protocol, which is why we reproduce Corollary 7.16 of [AH23] as Theorem 18 here.
The redeeming feature of the protocol is that the verifier only needs to perform spot checking very infrequently. The verifier enjoys significantly more time to perform these exponentially expensive calculations for each circuit. There are also modified protocols and modified adversary models tailored for practical implementation [L+25], which we do not discuss here.
8.1.1 Protocols
The above single round result allows one to use the entropy accumulation theorem to define a multi-round protocols that outputs certified smooth min-entropy for each case. Moreover, for randomness expansion, fresh randomness is only consumed on logarithmically many rounds to generate fresh challenge circuits.
We see that the random oracle model allows one to obtain a bound on the von Neumann entropy with quantum side information, and the respective RCS-based certified randomness protocols are certified randomness from entropy accumulation protocols satisfying Definition 8. We have omitted discussions of a general device in Section 5 of [AH23] since does not consider quantum side information, and Definition 8 is not satisfied. We have similarly omitted discussions of a semi-honest device in Section 6 of [AH23] since it only serves as inspiration for a more general adversary in Section 7.
For an instantiation based on Theorem 18, we use following protocol in Fig. 8, which uses the construction of the sequential compiler. However, the protocol in Fig. 8 is very different from the protocol in Fig. 4 of [AH23]. Specifically, we do not reuse the circuit and therefore do not perform the test by summing the scores over epochs. The main reason for this choice is that we do not believe the application of entropy accumulation in [AH23] is correct, which we discuss in the supplement Section 8.1.2.
From this protocol, the prover is given explicit description of the challenge circuits, not just oracle access, which is inevitable for protocols based on classical communication. Once again, given limitations on the prover’s computational power, the oracle model relies on the hope that the prover cannot do better than using the circuits as oracles even with explicit descriptions.
Input: the qubit count $n$, the number of rounds $\ell=\mathsf{poly}(\lambda)$, the score parameter $\delta\in[0,1]$ and the fraction of test rounds $\gamma=O((\log n)/\ell)$. Additionally, we require a quantum circuit ansatz over $n$ qubits defined as a function $C:\{0,1\}^{\mathsf{poly}(n)}\rightarrow U(n)$, and family of crytographic hash function $\{G_{k}\}_{k\in\{0,1\}^{\lambda}}:\{0,1\}^{m}\rightarrow\{0,1\}^{\mathsf{poly}(n)}$.
@@ 8.1.1 Protocols
  1. 1.
    At time $t=-\infty$, the verifiers sample a random hash key $k\leftarrow\{0,1\}^{\lambda}$. For $i\in[\ell]$, they sample random inputs $x_{i},y_{i}\in\{0,1\}^{m}$, a random challenge $\mathsf{ch}_{i}\in\{0,1\}^{\mathsf{poly}(n)}$, and $T_{i}\sim\mathsf{Bernoulli}(\gamma)$. They computes $s_{i}=G_{k}(x_{i}\oplus y_{i})\oplus\mathsf{ch}_{i}$ and $C_{i}=C(\mathsf{ch}_{i})$. They publish the hash key $k$.
  2. 2.
@@
    1. (a)
      At time $t=i-1$, $V_{0}$ sends $(x_{i},s_{i})$ and expects an answer $\mathsf{ans}_{i}$ at time $t=i$.
    2. (b)
      Similarly, at time $t=i-1$, $V_{1}$ sends $y_{i}$ and expects an answer $\mathsf{ans}^{\prime}_{i}$ at time $t=i$.
    3. (c)
      At time $t=i-1/2$, the honest prove, located at position $0.5$, computes $\mathsf{ch}_{i}=G_{k}(x_{i}\oplus y_{i})\oplus s_{i}$ and $C_{i}=C(\mathsf{ch}_{i})$, samples bitstrings $\mathsf{ans}_{i}=A_{i}$ from circuit $C_{i}$, and immediately sends $A_{i}$ to both verifiers.
@@
    $V$ accepts iff $\mathsf{ans}_{i}=\mathsf{ans}^{\prime}_{i}$ for all $i\in[\ell]$, and $\frac{1}{t}\sum_{i:T_{i}=1}p_{C_{i}}(A_{i})\geq(1+\delta)/N$ where $t=|\{i:T_{i}=1\}|$.
@@
We note that the original entropy accumulation theorem does not apply since Theorem 18 give entropy lower bounds for some scores, and the score is averaged over all rounds. This is in contract to bounding the entropy given some probability distribution. We discuss how to address this issue in the supplement Section 8.1.3.
8.1.2 Issues with the Protocol in Aaronson and Hung (STOC 23)
@@
One may argue that this type of argument could be used against certified randomness based on post-quantum secure trapdoor claw-free functions [BCM+21]. Indeed, [BCM+21] explicitly discusses this issue that entropy accumulation theorem requires that single-round entropy bound for all possible input states, including those that are computationally inefficient strings. As a result, [BCM+21] presents significant additional analysis to show that entropy accumulates in the protocol, and nontrivial work was presented in [MAF23] to use the entropy accumulation theorem.
@@
8.1.3 Issues with Entropy Accumulation
The single-round entropy lower bound in Theorem 18 is conditioned on the output achieving some score (e.g. $b$ in $b$-$\mathsf{XHOG}$), and the min-tradeoff functions used in [AH23] are defined for continuous scores. As a result, [AH23] developed a modified entropy accumulation theorem, which requires the Markov chain condition. However, for the arguments in Section 7, we prove that the modified adversary satisfies the non-signalling condition which allows us to apply generalized entropy accumulation. To apply the entropy accumulation theorem in [AH23] to the modified adversary, we have to show that the modified adversary also satisfies the Markov chain condition.
@@ 8.1.3 Issues with Entropy Accumulation
For the quantum channel defined in E.q. 89, for some $\mathcal{R}^{\prime}$, the Markov chain condition $A_{i}\leftrightarrow I_{i}\leftrightarrow I_{i+1}$ is satisfied:
@@
As a result, we can bound $H_{\rm min}^{\varepsilon}(A_{1}^{n}|C_{1}^{n}G_{1}^{n})$ by bounding $H_{\rm min}^{\varepsilon}(A_{1}^{n}|I_{1}^{n})$ instead with the entropy accumulation theorem of [AH23]. We can expand the definition of certified randomness from entropy accumulation to the case where the min-tradeoff function is for a score. To do this, we simply need to change $q$ for $\Sigma_{i}(q)$ into a score $s$, and define $h$ as $h=\min_{s}f(s)$. With this expanded definition, the protocols in Fig. 8 are certified randomness from repetition due to Theorem 18, and they are therefore CVPV protocols.
@@
8.2 Instantiation using Error-Correcting Codes
In this section, we show how to instantiate our single-round compiler (Theorem 7) using the seminal work of [YZ24], which results in a single-round CVPV scheme secure in QROM assuming the Aaronson-Ambainis conjecture (Theorem 20).
The following is taken verbatim from [YZ24], and is commonly known as the Aaronson-Ambainis Conjecture, originally due to [AA14]:
Conjecture 1 ([YZ24]).
Let $\varepsilon,\delta>0$. Given any quantum algorithm $\mathcal{A}$ that makes $Q$ quantum queries to a random oracle $H:\{0,1\}^{n}\to\{0,1\}^{m}$, there exists a deterministic classical algorithm $\mathcal{B}$ that makes $\mathsf{poly}(Q,m,\varepsilon^{-1},\delta^{-1})$ classical queries and satisfies
$\Pr_{H}\left[\left|\Pr\left[\mathcal{A}^{H}()\to 1\right]-\mathcal{B}^{H}()\right|\leq\varepsilon\right]\geq 1-\delta.$
8.2.1 Certified Randomness in QROM - Definitions
In [YZ24], single-round certified randomness as we define (Definition 3) is referred to as proof of min-entropy.
The quantum advantage of [YZ24] can be viewed as a quantum-query advantage, meaning they show a soundness gap between a quantum-query adversary and any bounded classical-query adversary. Given this definition, they show that any non-interactive quantum-query advantage gives certified randomness. Intuitively, a deterministic quantum algorithm would not have advantage over a classical-query algorithm, which can simulate a quantum-query algorithm assuming 1.
Definition 12 (Certified Randomness in QROM).
Let $G$ be a random oracle. A PoQ protocol $\mathscr{P}^{G}$ is said to have certified randomness property in the quantum random oracle model if no pair of a QPT prover $P^{G}$ and an unbounded guesser $Q^{G}$ can succeed in the following security game with non-negligible probability:
  1. 1.
    The verifier $V^{G}$ of $\mathscr{P}$ sends a challenge $\mathsf{ch}$ to both $P^{G}$ and $Q^{G}$.
  2. 2.
    $P^{G}$ sends back an answer $\mathsf{ans}$ and $Q$ and outputs a guess $\mathsf{ans}^{\prime}$.
  3. 3.
    $(P^{G},Q^{G})$ win if $V^{G}$ accepts and $\mathsf{ans}=\mathsf{ans}^{\prime}$.
We also mention the relevant definition from [YZ24] for completeness, presenting it in the operational (game-based) way we adopted for our other definitions.
Definition 13 (Proof of Min-Entropy in (auxiliary-input) QROM [YZ24]).
A (single-round) proof of min-entropy in the quantum random oracle model is a PoQ protocol $\mathscr{P}_{h}^{G}$, addittionally parametrized by a min-entropy threshold $h(\lambda)$, such that for any pair of a QPT prover $P^{G}$ and an unbounded guesser $Q^{G}$ the following is true: Consider the security game defined in Definition 12. For any inverse polynomial $\delta$, with overwhelming probability over $(G,\mathsf{ch})$ we have:
  • either the verifier $V^{G}_{h}$ accepts with probability at most $\delta$,
  • or the probability that $\mathsf{ans}=\mathsf{ans}^{\prime}$ conditioned on $V^{G}_{h}$ accepting is at most $2^{-h}$.
Remark 7 (Comparing Definitions).
Definition 12 can be thought of as the game-based version of the definition of [YZ24] in the auxiliary-input setting because we allow $Q^{G}$ to share an entangled state with $P^{G}$ that depends on $G$. Yet, Definition 13 is stronger than Definition 3 in several aspects. First, in Definition 13, the adversary has negligible success probability for an overwhelming fraction of challenges. Secondly, Definition 13 guarantees that the guesser $Q^{G}$ succeeds with exponentially small probability conditioned on $V^{G}$ accepting for an overwhelming fraction of $(G,\mathsf{ch})$, whereas Definition 12 gives a guarantee for average $(G,\mathsf{ch})$. Finally, in Definition 13 $P_{h}^{G}$ can output an answer with $h$-bits of min-entropy for any choice of polynomial $h(\lambda)$. We show the formal implication between the two definitions in the proof of Theorem 20.
8.2.2 Results
The following result was shown by prior work:
Theorem 19.
If 1 is true, then there exists a single-round proof of min-entropy in the quantum random oracle model.
This readily implies certified randomness with respect to Definition 12.
Corollary 2 (Certified Randomness in QROM [YZ24]).
If 1 is true, then there exists a certified randomness protocol in the quantum random oracle model.
@@ 8.2.2 Results
Let $\mathscr{P}^{G}$ be a protocol that satisfies Definition 3.5 in [YZ24]. We claim that it satisfies Definition 12 if we fix min-entropy as $h=\lambda$. We show a proof by contradiction.
Suppose there exists an adversary $(P^{G},Q^{G})$ that wins the security game with non-negligible probability. Let $E_{\mathsf{ACC}}$ be the event that $V^{G}$ accepts and $E_{\mathsf{GSS}}$ be the event that $\mathsf{ans}=\mathsf{ans}^{\prime}$. Similarly define $E_{\mathsf{ACC}}^{G,\mathsf{ch}}$ and $E_{\mathsf{GSS}}^{G,\mathsf{ch}}$ as the same events for fixed $G,\mathsf{ch}$. Then, for an inverse polynomial $\delta(\lambda)$ and an infinite set $\Lambda\subset\mathbb{Z}^{+}$ we have $\Pr\left[E_{\mathsf{ACC}}\land E_{\mathsf{GSS}}\right]\geq\delta$ for all security parameters $\lambda\in\Lambda$. By possibly shrinking $\Lambda$ we can assume that $\delta>2^{-\lambda}$ for $\lambda\in\Lambda$. For the rest of the proof, we restrict our attention to $\lambda\in\Lambda$.
By union bound, $\Pr\left[E_{\mathsf{ACC}}^{G,\mathsf{ch}}\land E_{\mathsf{GSS}}^{G,\mathsf{ch}}\right]\geq\delta/2$ for at least $\delta/2$ fraction of $(G,\mathsf{ch})$. In particular, $\Pr\left[E_{\mathsf{ACC}}^{G,\mathsf{ch}}\right]\geq\delta/2$ and $\Pr\left[E_{\mathsf{GSS}}^{G,\mathsf{ch}}|E_{\mathsf{ACC}}^{G,\mathsf{ch}}\right]\geq\delta/2>2^{-\lambda}$ for such $(G,\mathsf{ch})$. However, this contradicts with the guarantee of $\mathscr{P}^{G}$ because $\delta/2$ is non-negligible. ∎
Corollary 2 and Theorem 7 combined yield the following theorem:
Theorem 20 (Single-Round CVPV in QROM).
If 1 is true, then there exists a single-round CVPV protocol secure in the quantum random oracle model.
8.2.3 Construction
The reader may be curious to see what our single-round CVPV construction corresponding to Theorem 20 looks like when instantiated with real-world hash functions. We present it below at a high level for presentation purposes:
Construction 21.
Let $G_{k}$ and $F_{k^{\prime}}$ be cryptographic hash function families, and let $C$ be a suitably chosen error-correcting code.
  1. 1.
    The verifiers publish the hash key $k$, and set $s=G_{k}(x\oplus y)\oplus k^{\prime}$, where $x,y$ are random inputs.
  2. 2.
    At $t=0$, $V_{0}$ sends $(x,s)$ and $V_{1}$ sends $y$ to the prover simultaneously.
  3. 3.
    The honest prover, located at position $1$, computes $k^{\prime}=G_{k}(x\oplus y)\oplus s$. Then, he computes a codeword $\mathsf{ans}=(x_{1},x_{2},\dots,x_{m})\in C$ such that the $i$th bit of $F_{k^{\prime}}(x_{i})$ equals $1$ for $i\in[m]$. He immediately sends $\mathsf{ans}$ to both verifiers.
  4. 4.
  5. 5.
    The verifiers accept iff $\mathsf{ans}=\mathsf{ans}^{\prime}$ and $\mathsf{ans}=(x_{1},x_{2},\dots,x_{m})\in C$ such that the $i$th bit of $F_{k^{\prime}}(x_{i})$ equals $1$ for $i\in[m]$.
8.3 Instantiation using NTCFs
In this section, we show how to neatly relate the results of [BCM+21] and [LLQ22] using our work. First, we note that we can more explicitly use the result of [BCM+21] to get CVPV from LWE. Second, we note that there is an implication in the reverse direction.
8.3.1 CVPV from LWE in QROM
The result below is implicit in the work of [BCM+21], yet not formally stated before.
Lemma 17.
Assuming LWE, there exists a multi-round certified randomness scheme which satisfies Definition 10.
In fact, the stronger version with regular (not weak) certified randomness follow from the analysis of [BCM+21]. Now, we could recover the result of [LLQ22] using Theorem 14. Note that our compiler yields an entirely different (rapid-fire) construction.
Corollary 3 (CVPV from LWE).
If Lemma 17 is true, then there exists a secure CVPV scheme in the random oracle model.
8.3.2 (Weak) Certified Randomness from LWE using CVPV
Conversely, we can show Lemma 17 is true assuming the existence of a CVPV scheme secure under LWE. This could be of independent interest as it provides a formal argument as to the existence of an NTCF-based certified randomness protocol using our definition.
Proof of Lemma 17..
Assuming LWE, there exists a multi-round CVPV scheme secure in the random oracle model [^5] The results in the plain model require additional restrictions on the amount of entanglement shared between the adversaries. due to [LLQ22]. Using our compiler (Theorem 15), there exists a weak certified randomness protocol. Examining the compiler, we see that the resulting scheme uses the random oracle in a trivial way, which yields a weak certified randomness scheme in the plain model. ∎